OpenStudy (anonymous):

What does it mean to the rise over run when the slope is an integer? a. the rise number is one c. the run part of the slope is going to be one b. the run number is always negative d. there will be no slope

OpenStudy (freckles):

any integer can be written as a fraction

OpenStudy (freckles):

5 can be written as 5/1 -5 can be written as 5/1 0 can be written as 0/1

OpenStudy (freckles):

second line type-o -5 can be written as -5/1

OpenStudy (anonymous):

it is a the rise number is 1 because rise/run you allways put put the number over 1

OpenStudy (freckles):

@cool41 do you mean the run number is 1?

OpenStudy (freckles):

\[\frac{rise}{run}=integer=\frac{integer}{1} \]

OpenStudy (freckles):

thought you could also write it as \[\frac{-integer}{-1} \text{ since } \frac{-1}{-1}=+1\] so this problem is kinda wacky